S.J has 2 jobs offers at as a car salesperson. With the 1st one she will receive $600 per mth plus a 2% commission of the total sale of the automobiles she sells. At the other dealership, she will receive $1000 per mth plus a commission of 1.5% of her total sales.
34.which is the total sale of the automobiles that Ms. J must sell each mth to make the same income from either dealership?
35. Explain which job is better.

Answer by stanbon(75887)

Ms.J has 2 jobs offers as a car salesperson.
With the 1st one she will receive $600 per mth plus a 2% commission of the total sale of the automobiles she sells.
---
1st(x) = 600 + 0.02x
----------------------------------------
At the other dealership, she will receive $1000 per mth plus a commission of 1.5% of her total sales.
---
2nd(x) = 1000 + 0.015x
----------------------------------------

34.which is the total sales of the automobiles that Ms. J must sell each mth to make the same income from either dealership?
---
Solve: 1st(x) = 2nd(x)
600 + 0.02x = 1000 + 0.015x
0.005x = 400
x = 80,000 (break even point)
--------------------------------------------
35. Explain which job is better.
I'll leave that to you.
Compare the two programs based on x = 70,000
Which one provides the higher Income?
---
You could also graph the two plans on the same
coordinate system and draw a conclusion.
============================================
Cheers,
Stan H.